The product of two positive numbers is 105 and A.M is 11,then the largest number is
 

Grade:10

1 Answers

Arun
4 years ago
Dear Kavya
 
let numbers are a and b
 
Now ab = 105 and (a+b)/2 = 11
hence a + b = 22
 
Now (a – b)^2 = (a +b)^2 – 4ab = 484 – 420 = 64
 
hence a – b = 8
 
hence 2a = 30
a = 15 then b = 7